OpenStudy (anonymous):

What is this pattern 1,6,26,106,_,_,_ rule

OpenStudy (anonymous):

+5, +20, +80 See it?

OpenStudy (anonymous):

5(2^(2n))???

OpenStudy (anonymous):

\[a _{n+1}=a _{n}+5(2^{2n})?\]

OpenStudy (anonymous):

the 1st 1 being 1? thats the only thing i could come up with

OpenStudy (anonymous):

or 5(4^n)

OpenStudy (anonymous):

It's the difference X 4 then add the last number. 1,6,26,106,426,1706 1-6=5x4=20+6=26 6-26=20x4=80+26=106 26-106=80x4=320+106=426 106-426=320x4=1280+426=1706